im doing an assignment about equations, but i found one problem. my problem was found in this one:
FIND TWO POSITIVE NUMBERS WHOSE SUM IS 100 AND WHOSE PRODUCT IS AS LARGE AS POSSIBLE.
please help me, i must get my assignment done. thanx a lot! Sandra

Answer by checkley71(8403)

IF THEY CAN BE THE SAME NUMBER THEN THE ANSWER IS:
50+50=100
50*50=2,500
IF THEY HAVE TO BE DIFFERENT NUMBERS THEN THEY ARE:
49+51=100
49*51=2,499